# On putative self-similarity for incompressible 3D Euler

### v1

@@ 3.5. An outgoing property?
has $1$ as an eigenvalue, with associated eigenvector $\Omega(y_{*})$. Now $\mathbb{S}_{y_{*}}$ is a real, symmetric, traceless matrix ($\nabla\cdot U=0$) and we denote its eigenvalues (which are real) as $1,\lambda_{2},\lambda_{3}$, with $\lambda_{2}+\lambda_{3}=-1$.
If $\gamma y_{*}+U(y_{*})=0$, for $|y-y_{*}|\ll 1$ we write
$\gamma|y|^{2}+y\cdot U(y)$ $=\gamma|y-y_{*}|^{2}+(y-y_{*})^{i}(y-y_{*})^{j}(\partial_{j}U^{i})(y_{*})+\mathcal{O}(|y-y_{*}|^{3})$
$=\gamma|y-y_{*}|^{2}+(y-y_{*})^{i}(y-y_{*})^{j}\mathbb{S}_{y_{*}}^{ij}+\mathcal{O}(|y-y_{*}|^{3}).$
Thus, if the outgoing condition (3.34) holds in an open neighborhood of $y_{*}$, then for all $|y-y_{*}|\ll 1$ we have $\widehat{(y-y_{*})}\cdot\mathbb{S}_{y_{*}}\widehat{(y-y_{*})}\geq c_{*}-\gamma+\mathcal{O}(|y-y_{*}|)$. Passing $y\to y_{*}$, we deduce that the smallest eigenvalue of the matrix $\mathbb{S}_{y_{*}}$, satisfies $\lambda_{\rm min}(\mathbb{S}_{y_{*}})\geq c_{*}-\gamma$. Thus, the sum of the two-smallest eigenvalues of $\mathbb{S}_{y_{*}}$ is $\geq 2(c_{*}-\gamma)$, and due to the zero trace property, the largest eigenvalue of the matrix $\mathbb{S}_{y_{*}}$ satisfies $\lambda_{\rm max}(\mathbb{S}_{y_{*}})\leq 2(\gamma-c_{*})$. We have thus shown that all eigenvalues of $\mathbb{S}_{0}$ lie in the interval $\lambda_{\rm min}(\mathbb{S}_{y_{*}}),\lambda_{\rm max}(\mathbb{S}_{y_{*}})]\subseteq[c_{*}-\gamma,2(\gamma-c_{*})]$. In particular, because the eigenvalue $1$ lies in this interval, we deduce $1\leq 2(\gamma-c_{*})$. The claimed lower bound now follows. ∎
@@
Let $U$ be a $C^{2}$ smooth similarity profile solving (3.3). Let $y_{*}\in\mathbb{R}^{3}$ be a zero of $\gamma y+U(y)$. [^7] Such a zero of $\gamma y+U(y)$ always exists: $y=0$. If $U$ satisfies the outgoing property (3.34) locally near $y=y_{*}$ for some $c_{*}\geq 0$, and if $\gamma<1/2+c_{*}$, then we must have $\Omega(y_{*})=0$. Additionally, for any $n\in\mathbb{N}$ such that $\Omega$ is $C^{n+1}$ smooth in an open neighborhood of $y_{*}$, we have that $\nabla^{n}\Omega(y_{*})=0$. That is, when $\gamma<1/2+c_{*}$, a smooth vorticity profile must vanish to infinite order at $y=y_{*}$.
@@
Let $\Omega$ be a nontrivial $C^{1}$ smooth globally self-similar vorticity profile for 3D incompressible Euler equations. Assume that the outgoing property (3.34) holds globally for all $y\in\mathbb{R}^{3}$ for some $c_{*}\geq 0$. We distinguish two cases:
  • (i)
    If $c_{*}>0$ and if $\Omega$ is real-analytic at $y=0$ (has a convergent power series at the origin), then
    $\gamma\geq\frac{1}{2}+c_{*}.$
  • (ii)
    If $c_{*}=0$, if we assume that all zeros $y_{*}$ of $\gamma y+U(y)$ are isolated and that $\Omega$ is real-analytic [^8] The high regularity of the profile in a neighborhood of fixed points of the self-similar Lagrangian flow played a crucial role in the proofs of finite time blowup for the compressible Navier-Stokes equations via self-similar imploding solutions of compressible Euler; see the arguments presented in [36, 3, 41]. These papers strongly suggest that lifting a self-similar singularity for 3D incompressible Euler to 3D incompressible Navier-Stokes, would also require this infinite local regularity at Lagrangian fixed points. at each of these points $y_{*}$, then
@@
We note that while the conclusion of Theorem 3.9 is the same as that of Theorem 3.7, the assumptions are different. Here we do not assume $\Omega(y_{*})\neq 0$, and instead work with the complementary case when $\Omega(y_{*})=0$; the extra assumptions are infinite regularity at the sonic point $y_{*}$ (it is reasonable to expect that for $|y|\ll 1$,  (3.4) could be solved via a power series expansion), and that the outgoing property holds globally, not just near the origin, for some $c_{*}\geq 0$.
@@
Let us first give the proof in the simpler case $c_{*}>0$. Assume by contradiction that $\gamma<1/2+c_{*}$. Due to Proposition 3.8 we know that $\nabla^{n}\Omega(0)=0$ for all $n\in\mathbb{N}_{0}$. The analyticity assumption at $y=0$ then implies the existence of a radius (of analyticity) $R_{*}>0$ such that $\Omega(y)=0$ for all $|y|<R_{*}$.
The global outgoing property with $c_{*}>0$ implies that $\frac{1}{2}\frac{d}{d\tau}|Y(a,\tau)|^{2}\geq c_{*}|Y(a,\tau)|^{2}$, and thus $|Y(a,\tau)|\geq|a|e^{c_{*}\tau}$, for any label $a\in\mathbb{R}^{3}\setminus\{0\}$. Thus, the back-to-labels map $Y^{-1}(\cdot,\tau)\colon\mathbb{R}^{3}\to\mathbb{R}^{3}$ satisfies
$|Y^{-1}(y,\tau)|\leq|y|e^{-c_{*}\tau},\qquad\mbox{for any}\qquad|y|>0.$
@@
Therefore, for any given $y\in\mathbb{R}^{3}$ with $|y|\geq R_{*}$, there exists a finite time $\tau_{*}=\tau_{*}(|y|,c_{*},R_{*})\geq 0$, such that $|{Y^{-1}}(y,\tau)|<R_{*}$ for all $\tau>\tau_{*}$.
At this point we appeal to the self-similar Cauchy formula (3.22), which states that
$\Omega(y)=e^{-(\gamma+1)\tau}(\nabla_{a}Y)(Y^{-1}(y,\tau),\tau)\Omega(Y^{-1}(y,\tau)).$
For any $|y|\in\mathbb{R}^{3}$ with $|y|\geq R_{*}$, once we evaluate the above identity at a time $\tau>\tau_{*}(|y|,c_{*},R_{*})\geq 0$, and using that $\Omega\equiv 0$ on $B_{R_{*}}(0)$, we deduce $\Omega(y)=0$. Thus $\gamma<1/2+c_{*}$ leads to a contradiction.
Next, we consider the degenerate case $c_{*}=0$. For some integer $N\geq 0$, let $\{y_{*}^{(\ell)}\}_{\ell=0}^{N}\in\mathbb{R}^{3}$ be all the zeros of $\gamma y+U(y)$. [^9] There can be at most finitely many such zeros $\gamma y+U(y)$ because all these zeros must lie in the ball of radius $(\frac{2\mathsf{C}_{\flat}}{\gamma})^{\gamma}$ near the origin. If infinitely many zeros would exist, they would need to have an accumulation point, contradicting the assumption that the zeros of $\gamma y+U(y)$ are isolated. Our convention is that $y_{*}^{(0)}=0$. If $\gamma<1/2$, by Proposition 3.8 we have $\nabla^{n}\Omega(y_{*}^{(\ell)})=0$ for all $n\in\mathbb{N}_{0}$, and then analyticity of $\Omega$ near $y_{*}^{(\ell)}$ shows the existence of an $R_{*}^{(\ell)}>0$ such that $\Omega(y)=0$ for all $|y-y_{*}^{(\ell)}|<R_{*}^{(\ell)}$, for all $0\leq\ell\leq N$.
In order to complete the proof, via the self-similar Cauchy formula (3.22), we need to show that for all $y\in\mathbb{R}^{3}$ there exists a finite time $\tau_{*}\geq 0$ such that
$Y^{-1}(y,\tau_{*})\in\mathcal{D}:=\bigcup\limits_{\ell=0}^{N}B_{R_{*}^{(\ell)}}\bigl(y_{*}^{(\ell)}\bigr).$ (3.37)
@@
Due to the discussion in the first paragraph of Section 3.5, we only need to check condition (3.37) for $y$ belonging to the compact set $\mathcal{D}^{\complement}\cap\overline{B_{R_{\flat}}(0)}$, where $R_{\flat}:=(\frac{2\mathsf{C}_{\flat}}{\gamma})^{\gamma}$. [^10] This is because when $|y|>R_{\flat}$, we have that $|Y^{-1}(y,\tau)|\leq|y|e^{-\gamma\tau/2}$, as long as $|Y^{-1}(y,\tau)|\geq R_{\flat}$. By the continuity of $U$, and since all the zeros of $\gamma y+U(y)$ lie in the interior of $\mathcal{D}$, there exists a $c_{\flat}>0$ such that
$y\cdot\bigl(\gamma y+U(y)\bigr)\geq c_{\flat}|y|^{2},\qquad\mbox{for all}\qquad y\in\mathcal{D}^{\complement}\cap\overline{B_{R_{\flat}}(0)}.$
@@
Moreover, recall that if $y\in\overline{B_{R_{\flat}}(0)}$, then for all $\tau>0$ we have that $Y^{-1}(y,\tau)\in B_{R_{\flat}}(0)$. Combining these to facts shows that for any $y\in\mathcal{D}^{\complement}\cap\overline{B_{R_{\flat}}(0)}$, for all $\tau>0$ such that $Y^{-1}(y,\tau)\not\in\mathcal{D}$, we have $|Y^{-1}(y,\tau)|\leq|y|e^{-c_{\flat}\tau}$. As $\mathcal{D}$ contains an open ball near the origin, there exists a finite time $\tau_{*}>0$ such that $Y^{-1}(y,\tau)\in\mathcal{D}$, thereby completing the proof. ∎
@@ 4.2. Consequences of an axisymmetric outgoing property
We revisit the discussion from Section 3.5, in the axisymmetric setting. The outgoing assumption (3.34) becomes
$\gamma(r^{2}+z^{2})+rU_{r}(r,z)+zU_{z}(r,z)\geq c_{*}(r^{2}+z^{2})$ (4.6)
for some $c_{*}\geq 0$. When $c_{*}>0$, the above condition implies that the only fixed point for the ODE (4.4) is $(r,z)=(0,0)$. When $c_{*}=0$, other fixed points away from the origin may exist.
In this section we only record the results of Theorems 3.7 and 3.9 in the axisymmetric setting. Sections 4.3 and 4.4 contain the new results: therein we analyze the case when the outgoing property (3.34) fails.
@@
    If $\Omega_{z}(0,0)\neq 0$ and if the outgoing property (4.6) holds locally near $(r,z)=(0,0)$ for some $c_{*}\geq 0$, then $\gamma\geq 1/2+c_{*}$.
@@
    If $\Omega_{z}(0,0)=0$, if $\Omega$ is real analytic at $(r,z)=(0,0)$, and if the outgoing property (4.6) holds globally for all $(r,z)\in\mathbb{R}_{+}\times\mathbb{R}$ for some $c_{*}>0$, then $\gamma\geq 1/2+c_{*}$.
  • (iii)
    If the outgoing property (4.6) holds globally for all $(r,z)\in\mathbb{R}_{+}\times\mathbb{R}$ with $c_{*}=0$, if all the fixed points of the ODE system (4.4) are isolated, and if $\Omega$ is real-analytic at each of these points, then $\gamma\geq 1/2$.
@@ 4.3. Failure of the outgoing property: fixed points with nonzero swirl velocity
Here we explore constraints imposed by the existence of a fixed point $(r,z)\in\mathbb{R}_{+}\times\mathbb{R}$ for the (4.4) dynamics; i.e., points such that
@@
In this case, we automatically deduce $R(r,z,\tau)=r$, $Z(r,z,\tau)=z$, for all $\tau\geq 0$, and $\Theta(r,z,\tau)=\frac{\tau}{r}U_{\theta}(r,z)\mod 2\pi$. If we additionally know that $U_{\theta}(r,z)\neq 0$, [^11] The condition $U_{\theta}(r,z)\neq 0$ implies $r>0$, due to Remark 4.2. we have thus obtained a periodic orbit for the full dynamics of the Lagrangian flow map $Y(\cdot,\tau)$ with nonzero circulation. This leads to the following result:
@@
If $(r,z)$ is a solution of (4.7) with $U_{\theta}(r,z)=0$, then also $\Omega_{\theta}(r,z)=0$. This follows from the swirl component of the vorticity equation (4.3) and from incompressibility.
@@
$e^{(1-2\gamma)\tau}\Gamma_{\rm ss}(\tau)=e^{(1-2\gamma)\tau}\int_{\mathbb{T}}U_{\theta}(r,z)rd\lambda=e^{(1-2\gamma)\tau}\cdot 2\pi rU_{\theta}(r,z).$

### v2

@@ 3.5. An outgoing property?
has $1$ as an eigenvalue, with associated eigenvector $\Omega(y_{*})\neq 0$. Now $\mathbb{S}_{y_{*}}$ is a real, symmetric, traceless matrix ($\nabla\cdot U=0$) and we denote its eigenvalues (which are real) as $1,\lambda_{2},\lambda_{3}$, with $\lambda_{2}+\lambda_{3}=-1$.
Since $V(y_{*})=0$, for $|y-y_{*}|\ll 1$ we have
$V(y)\cdot(y-y_{*})$ $=(y-y_{*})\otimes(y-y_{*})\colon\!(\nabla V)(y_{*})+\mathcal{O}(|y-y_{*}|^{3})$
$=(y-y_{*})\otimes(y-y_{*})\colon\!(\gamma\mathrm{Id}+\mathbb{S}_{y_{*}})+\mathcal{O}(|y-y_{*}|^{3}).$
Thus, if the outgoing condition (3.37) holds in an open neighborhood of $y_{*}$, then for all $|y-y_{*}|\ll 1$ we have $\gamma+\hat{z}\cdot\mathbb{S}_{y_{*}}\hat{z}+\mathcal{O}(|y-y_{*}|)\geq c_{*}$, where $\hat{z}=\frac{y-y_{*}}{|y-y_{*}|}$. Passing $y\to y_{*}$, we deduce that the smallest eigenvalue of the matrix $\mathbb{S}_{y_{*}}$, satisfies $\lambda_{\rm min}(\mathbb{S}_{y_{*}})\geq c_{*}-\gamma$. Thus, the sum of the two-smallest eigenvalues of $\mathbb{S}_{y_{*}}$ is $\geq 2(c_{*}-\gamma)$, and due to the zero trace property, the largest eigenvalue of the matrix $\mathbb{S}_{y_{*}}$ satisfies $\lambda_{\rm max}(\mathbb{S}_{y_{*}})\leq 2(\gamma-c_{*})$. We have thus shown that the eigenvalues of $\mathbb{S}_{y_{*}}$ satisfy
$\lambda_{\rm min}(\mathbb{S}_{y_{*}}),\lambda_{\rm max}(\mathbb{S}_{y_{*}})]\subseteq[c_{*}-\gamma,2(\gamma-c_{*})].$
In particular, because the eigenvalue $1$ lies in this interval, we deduce $1\leq 2(\gamma-c_{*})$. The claimed lower bound $\gamma\geq c_{*}+1/2$ now follows.∎
@@
Let $U$ be a $C^{2}$ smooth similarity profile solving (3.3). Let $y_{*}\in\mathbb{R}^{3}$ be a zero of $V(y)=\gamma y+U(y)$. If $U$ satisfies the outgoing property (3.37) locally near $y=y_{*}$ for some $c_{*}\geq 0$, and if $\gamma<1/2+c_{*}$, then we must have $\Omega(y_{*})=0$. Additionally, for any $n\in\mathbb{N}$ such that $\Omega$ is $C^{n+1}$ smooth in an open neighborhood of $y_{*}$, we have that $\nabla^{n}\Omega(y_{*})=0$. That is, when $\gamma<1/2+c_{*}$, a smooth vorticity profile must vanish to infinite order at $y=y_{*}$.
@@
Let $U$ be a nontrivial $C^{2}$ smooth globally self-similar velocity profile for 3D incompressible Euler equations. Assume that the local outgoing property of Definition 3.7 holds. Furthermore, assume that $\Omega$ is real-analytic at all $y_{*}\in\mathcal{N}_{V}$. [^7] The high regularity of the profile in a neighborhood of fixed points of the self-similar Lagrangian flow played a crucial role in the proofs of finite time blowup for the compressible Navier-Stokes equations via self-similar imploding solutions of compressible Euler; see the arguments presented in [35, 3, 41]. These papers strongly suggest that lifting a self-similar singularity for 3D incompressible Euler to 3D incompressible Navier-Stokes, would also require this infinite local regularity at Lagrangian fixed points. Then,
@@
We note that while the conclusion of Theorem 3.10 is similar to that of Theorem 3.8, the assumptions are different. Here we do not assume $\Omega(y_{*})\neq 0$, and instead work with the complementary case when $\Omega(y_{*})=0$; the extra assumption is infinite regularity at the nodal points $y_{*}$ of $V$ (it is reasonable to expect that for $|y-y_{*}|\ll 1$, (3.4) could be solved via a power series expansion).
@@
Assume by contradiction that $\gamma<1/2$. For some integer $N\geq 0$, let $\mathcal{N}_{V}=\{y_{*}^{(\ell)}\}_{\ell=0}^{N}\in\mathbb{R}^{3}$. Our convention is that $y_{*}^{(0)}=0$. Since $\gamma<1/2+c_{*}$, by Proposition 3.9 we have $\nabla^{n}\Omega(y_{*}^{(\ell)})=0$ for all $n\in\mathbb{N}_{0}$. The analyticity of $\Omega$ near $y_{*}^{(\ell)}$ shows the existence of a radius $R_{*}>0$ such that $\Omega(y)=0$ for all $|y-y_{*}^{(\ell)}|<R_{*}$, for all $0\leq\ell\leq N$.
@@
Next, we claim that for all $y_{0}\in\mathbb{R}^{3}$ there exists a finite time $\tau_{0}=\tau_{0}(y_{0})\leq 0$ such that the backwards-in-time flow map satisfies
$Y(y_{0},\tau_{0})\in\mathcal{D}:=\bigcup\limits_{\ell=0}^{N}B_{R_{*}}\bigl(y_{*}^{(\ell)}\bigr).$ (3.40)
@@
Due to the discussion in the first paragraph of Section 3.5, we only need to check condition (3.40) for $y_{0}$ belonging to the compact set $\mathcal{K}:=\mathcal{D}^{\complement}\cap\overline{B_{R_{\flat}}(0)}$, where $R_{\flat}$ is as defined in (3.35). [^8] This is because when $|y_{0}|>R_{\flat}$, we have that $|Y(y_{0},\tau)|\leq|y_{0}|e^{\gamma\tau/2}$, as long as $|Y(y_{0},\tau)|\geq R_{\flat}$. Thus, there exists a sufficiently negative $\tau$ such that $|Y(y_{0},\tau)|=R_{\flat}$. Since all the zeros of $V(y)=\gamma y+U(y)$ lie in the interior of $\mathcal{D}$, since $V$ is continuous and $\mathcal{K}$ is compact, there exists a $c_{\flat}>0$ such that
$|V(y)|\geq c_{\flat},\qquad\mbox{for all}\qquad y\in\mathcal{K}=\mathcal{D}^{\complement}\cap\overline{B_{R_{\flat}}(0)}.$
@@
Moreover, recall that if $y_{0}\in\overline{B_{R_{\flat}}(0)}$, then for all $\tau<0$ we have that $Y(y_{0},\tau)\in B_{R_{\flat}}(0)$. Combining these two facts shows that for any $y_{0}\in\mathcal{K}$ and for all $\tau<0$ such that $Y(y_{0},\tau)\not\in\mathcal{D}$, we have $|V(Y(y_{0},\tau))|\geq c_{\flat}$. At this stage we recall from (3.32) that the Bernoulli function $\mathcal{H}$ satisfies
$\frac{d}{d\tau}\mathcal{H}(Y(y_{0},\tau))=(2\gamma-1)|V(Y(y_{0},\tau))|^{2}\leq(2\gamma-1)c_{\flat}^{2}$
for all $\tau<0$ such that $Y(y_{0},\tau)\not\in\mathcal{D}$; here we have used again that $\gamma<1/2$. Integrating this inequality on $\tau,0]$ implies
$\mathcal{H}(Y(y_{0},\tau))\geq\mathcal{H}(y_{0})-\tau(1-2\gamma)c_{\flat}^{2}$
as long as $Y(y_{0},\tau)\in\mathcal{K}$. Since the Bernoulli function $\mathcal{H}$ is continuous and $\mathcal{K}$ is compact, we deduce that $\tau$ cannot be taken to be arbitrarily negative; there must exist $\tau_{0}\leq 0$ such that $Y(y_{0},\tau_{0})\in\mathcal{D}$. Since we have previously shown that $\Omega\equiv 0$ on $\mathcal{D}$, it follows that for any $y_{0}\in\mathbb{R}^{3}$ there exists $\tau_{0}\leq 0$ such that $\Omega(Y(y_{0},\tau_{0}))=0$.
To conclude the proof, we appeal to the self-similar Cauchy formula (3.22), which states that
$e^{(\gamma+1)\tau}\Omega(Y(y_{0},\tau))=(\nabla_{a}Y)(y_{0},\tau)\Omega(y_{0})$
for all $y_{0}\in\mathbb{R}^{3}$ and all $\tau\in\mathbb{R}$. Evaluating the above expression at $\tau=\tau_{0}$, using that $\Omega(Y(y_{0},\tau_{0}))=0$, and noting that $\mathrm{det}(\nabla_{a}Y)(y_{0},\tau_{0})=e^{3\gamma\tau_{0}}\neq 0$ (so that this matrix is invertible), we deduce $\Omega(y_{0})=0$. Since $y_{0}$ was arbitrary we deduce $\Omega\equiv 0$, a contradiction. Thus $\gamma\geq 1/2$. ∎
@@ 4.2. Consequences of an axisymmetric outgoing property
We revisit the discussion from Section 3.5, in the axisymmetric setting.
In axisymmetry, we have that
$V=\gamma y+U=\underbrace{(\gamma r+U_{r})}_{=:V_{r}}\vec{e}_{r}+U_{\theta}\vec{e}_{\theta}+\underbrace{(\gamma z+U_{z})}_{=:V_{z}}\vec{e}_{z}.$
Thus, $(r_{*},z_{*})$ is a zero of $V$ if and only if $\gamma r_{*}+U_{r}(r_{*},z_{*})=U_{\theta}(r_{*},z_{*})=\gamma z_{*}+U_{z}(r_{*},z_{*})=0$. For such a zero we note however that back in Cartesian $\mathbb{R}^{3}$, for any point $y_{*}$ on the circle $(r_{*},\theta,z_{*})_{\theta\in[0,2\pi]}$ we have that $V(y_{*})=0$ (by axisymmetry). Thus, if we insist on the finiteness of the nodal set $\mathcal{N}_{V}\subset\mathbb{R}^{3}$, as required by Definition 3.7, it follows that $r_{*}=0$; the aforementioned circle contracts onto a point. Hence the finiteness of $\mathcal{N}_{V}$ implies that all nodal values of $V$ must lie on the axis of symmetry, and hence with Remark 4.2,
$\mathcal{N}_{V}=\{(0,0,z_{j})\colon V_{z}(0,z_{j})=\gamma z_{j}+U_{z}(0,z_{j})=0,\;-M\leq j\leq N\}$
for some integers $M,N\geq 0$. For convenience we let $z_{0}=0$, which always yields a nodal value.
In the vicinity of zeroes $(0,z_{*})$ of $V$, inequality (3.37) from Definition 3.7 becomes
$rV_{r}(r,z)+(z-z_{*})V_{z}(r,z)\geq c_{*}r^{2}+c_{*}(z-z_{*})^{2},$ (4.6)
for all $(r,z)\in[0,\infty)\times\mathbb{R}$ with $r^{2}+(z-z_{*})^{2}\leq\varepsilon_{*}^{2}$.
We record the results of Theorems 3.8 and 3.10 in the axisymmetric setting. Sections 4.3 and 4.4 contain the new results: therein we analyze the case when the outgoing property of Definition 3.7 fails.
@@
    Let $(0,z_{*})$ be a zero of $V$, meaning that $\gamma z_{*}+U_{z}(0,z_{*})=0$. If $\Omega_{z}(0,z_{*})\neq 0$ and if the outgoing property (4.6) holds locally near $(0,z_{*})$ for some $c_{*}\geq 0$, then $\gamma\geq 1/2+c_{*}$.
@@
    If the nodal set $\mathcal{N}_{V}\subset\mathbb{R}^{3}$ is finite, the outgoing property (4.6) holds locally in the vicinity of all nodal values $(0,z_{*})\in\mathbb{R}_{+}\times\mathbb{R}$ for some $c_{*}\geq 0$, and if $\Omega$ is real-analytic at each of these nodal values, then $\gamma\geq 1/2$.
@@ 4.3. Failure of the outgoing property: fixed points with nonzero swirl velocity
Here we explore constraints imposed by the existence of a fixed point $(r,z)\in\mathbb{R}_{+}\times\mathbb{R}$ for the dynamics (4.4) in the meridional plane; i.e., points such that
@@
In this case, we automatically deduce $R(r,z,\tau)=r$, $Z(r,z,\tau)=z$, for all $\tau\in\mathbb{R}$, and $\Theta(r,z,\tau)=\frac{\tau}{r}U_{\theta}(r,z)\mod 2\pi$. If we additionally know that $U_{\theta}(r,z)\neq 0$, [^9] The condition $U_{\theta}(r,z)\neq 0$ implies $r>0$, due to Remark 4.2. we have thus obtained a periodic orbit for the full dynamics of the Lagrangian flow map $Y(\cdot,\tau)$ with nonzero circulation. This leads to the following result:
@@
If $(r,z)$ is a solution of (4.7) with $U_{\theta}(r,z)=0$, then also $\Omega_{\theta}(r,z)=0$. This follows from the swirl component of the vorticity equation (4.3).
@@
$\Gamma_{\rm ss}(0)=e^{(1-2\gamma)\tau}\Gamma_{\rm ss}(\tau)=e^{(1-2\gamma)\tau}\int_{\mathbb{T}}U_{\theta}(r,z)rd\lambda=e^{(1-2\gamma)\tau}\cdot 2\pi rU_{\theta}(r,z).$